### v2

# Arakelov geometry on flag varieties over function fields and related topics

Abstract.
Let $k$ be an algebraically closed field of characteristic zero. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
We compute the height filtration, successive minima, and the Boucksom-Chen concave transform of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$. An interesting application is that the height of $X$ equals to a weighted average of successive minima, and one may view this as a refinement of Zhang’s inequality of successive minima.
Let $f\in N^{1}(F/P)$ be the numerical class of a vertical fiber. We compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$, and it turns out that they are almost the same as the height filtration. As a corollary, we compute the $k$-th movable cones of flag bundles over curves for all $k$.
1. Introduction
1.1. Height filtration and successive minima
We start by giving a reminder on height functions in Arakelov theory. Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [24, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give the definition.
If $K=k(C)$, consider a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$ with $\mathcal{L}_{K}\simeq L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
$h_{\overline{L}}:X(\overline{K})\longrightarrow\mathbb{Q},\;x\longmapsto\frac{\mathcal{L}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}\;\;\text{where $\overline{\{x\}}$ is the closure of $x$ in $\mathcal{X}$}.$
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly via arithmetic intersection theory.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points the successive minima. Note that our definition of successive minima are slightly different with Zhang [25]. Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In this article, $e_{i}$ will be called the successive minima of Zhang to avoid ambiguity.
Determining the height filtration is generally difficult and only two cases are known in the literature:
  • The Néron-Tate height function on abelian varieties $A$ over $K$. Let $L$ be an ample symmetric line bundle over $A$ endowed with the unique adelic metric such that $[2]^{*}\overline{L}=2^{\dim A}\overline{L}$. The height filtration of the induced Néron-Tate height function $\widehat{h}_{L}$ is $A\supsetneq\emptyset$ and the jumping point is $\zeta=0$;
  • The toric height function on toric varieties [8]. They are given by successively deleting torus orbits. An important example is the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$ and the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$. The height filtration on $\mathbb{P}^{n}_{\mathbb{Q}}$ is given by
    $\quad\quad\quad\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\emptyset$
    with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
Our first result (Theorem A below) provides a new case where height filtrations can be explicitly computed. Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined.
Theorem A (Theorem 2.1).
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
In particular, the filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite (while in general height filtrations could be infinite) and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$, $w\in W_{Q}\backslash W/W_{P}$.
1.2. String polytopes and the Boucksom-Chen concave transform
Attached to a projective variety $X$ over $K$ and an adelic line bundle $\overline{L}$ on $X$ with $L$ ample and $\overline{L}$ semipositive, we have the Okounkov body $\Delta$ and the Boucksom-Chen concave transform $\varphi:\Delta\longrightarrow\mathbb{R}$ thanks to [17, 18, 6, 1]. The maximum of $\varphi$ equals to the essential minimum of $h_{\overline{L}}$, the minimum of $\varphi$ equals to the absolute minimum of $h_{\overline{L}}$ and the integral mean of $\varphi$ equals to the height of $X$.
In toric cases, they can be expressed as summation of local roof functions [7]. But in general, the Boucksom-Chen concave transform is difficult to compute. Our second result (Theorem B) provides a new case where the Boucksom-Chen concave transform can be explicitly computed.
To state the result, keep the notations in Theorem A. Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. Then we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$ [19, 3], which is a polytope of dimension $d=\dim G/P$. Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map. The Okounkov body of $(X,L)$ (with respect to a specific valuation) is the string polytope $\Delta_{\lambda}$.
Theorem B (Theorem 3.1).
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. As we mentioned above, the variety height is given by the integral mean of the Boucksom-Chen concave transform [6]. Thus a direct corollary is
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
The integration can be computed via Weyl symmetry.
Theorem C (Theorem 3.2).
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ in Theorem C are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$. Thus Theorem C builds an equality between the variety height and a weighted average of the successve mimina. One may recognize this as a refinement of Zhang’s inequality of successive minima [25]
$e_{1}\geq h_{\mathcal{L}_{\lambda}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
into an equality. Here $e_{i}$ is the the successive minima of Zhang.
We remark that it is not possible to build such an equality in general. Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$. The variety height is a rational number $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
1.3. Base loci and movable cones
It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ is equivalent to the positivity of the induced height function $h_{\overline{L}}$. The study of essential minimum and pseudo-effectivity [8, 1, 22] provides an evidence.
Our third result (Theorem D below) provides a new evidence to support the above mentioned philosophy. Keep the notations in Theorem A. Let $f\in N^{1}(F/P)$ be the class of a vertical fiber and $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$ for $t\in\mathbb{R}$. They measure the positivity of the line bundle $\mathcal{L}_{\lambda}$.
Theorem D (Theorem 4.1).
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
The height filtration $Z_{t}$ in Theorem A measures the positivity of the height function $h_{\mathcal{L}_{\lambda}}$, and we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is just the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
A class $\alpha\in N^{1}(F/P)_{\mathbb{R}}$ is called $k$-movable if $\mathrm{B}_{+}(\alpha)$ has codimension $\geq k$. As a direct corollary of Theorem D, we compute the movable cones on flag bundles. (Please see §4.3 for the notations.)
Theorem E (Theorem 4.2).
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ coincides with the cone defined by
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
Let $E$ be a vector bundle on a curve $C$. The ample cone of $\mathbb{P}(E)$ is computed in [21] and the big cone of $\mathbb{P}(E)$ is computed in [12]. Fulger and Lehmann computed $\operatorname{Mov}^{2}(\mathbb{P}(E))$ in [13]. In [4], the big cone of the Grassmann bundle of $E$ is computed and in [5], the ample cone of flag bundles of $E$ is computed. Our Theorem E is a simultaneous generalization of all these results.
Theorem F (Corollary 4.3).
Let $\mathcal{L}$ be a line bundle on $F/P$. Then $\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$. Here $e_{k}(h_{\mathcal{L}})$ is the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$.
Take $k=1$, we recover [1, Theorem 1.1] (in this special case) that $\mathcal{L}_{\lambda}$ big $\Longleftrightarrow$ $\zeta_{\operatorname{ess}}(h_{\mathcal{L}_{\lambda}})>0$.
1.4. Example: Grassmann bundles
We provide here the computation of the essential minimum and the big cone on Grassmann bundles.
Let $k$ be an algebraically closed field and $C$ be a curve over $k$ with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$ with Harder-Narasimhan filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $B$ be the subgroup of $\operatorname{GL}_{n}$ of all upper triangular matrices, and let $Q$ be the parabolic subgroup containing $B$ of type $(\operatorname{rank}(E_{1}),\operatorname{rank}(E_{2}),\dots,\operatorname{rank}(E_{r}))$. Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block trianguluar matrices of the shape
$\left[\begin{array}[]{cccc}A_{1}&*&\cdots&*\\ &A_{2}&\cdots&*\\ &&\cdots&*\\ &&&A_{r}\end{array}\right],\quad A_{i}\in\operatorname{GL}_{a_{i}}.$
Let $F(E)$ be the frame bundle of $E$, which is a $\operatorname{GL}_{n}$-principal bundle. Then the canonical reduction of $F(E)$ to $Q$ is $F_{Q}$, the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $\mu_{i}$ ($1\leq i\leq n$) be the $i$-th number in the sequence
$\underbrace{\mu(E_{1}),\dots,\mu(E_{1})}_{\text{$\operatorname{rank}(E_{1})$ times}},\underbrace{\mu(E_{2}/E_{1}),\dots,\mu(E_{2}/E_{1})}_{\text{$\operatorname{rank}(E_{2}/E_{1})$ times}},\dots,\underbrace{\mu(E/E_{r-1}),\dots,\mu(E/E_{r-1})}_{\text{$\operatorname{rank}(E/E_{r-1})$ times}}.$
Note that $X(T)_{\mathbb{Q}}=\bigoplus_{i=1}^{n}\mathbb{Q}\lambda_{i}$ where $\lambda_{i}:T\longrightarrow\mathbb{G}_{m}$ is the character of taking the $i$-th entry, and $\langle{\mathrm{deg}}(F_{Q}),\lambda_{i}\rangle=\mu_{i}$.
Let $\operatorname{Gr}_{r}(E)=F(E)/P$ be the Grassmann bundle of $r$-dimension quotients, where $P$ is the parabolic subgroup containing $B$ of type $(n-r,r)$. Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual and $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
$\left[\begin{matrix}1&2&\cdots&n-r&n-r+1&\cdots&n\\ *&*&\cdots&*&i_{1}&\cdots&i_{r}\end{matrix}\right]\longleftrightarrow(i_{1},\dots,i_{r}).$
$N^{1}(\operatorname{Gr}_{r}(E))$ is two dimensional with basis $\mathcal{O}(1)$ (determinant bundle of the universal bundle on $\operatorname{Gr}_{r}(E)$) and $f$ (a vertical fiber). Let $\det_{2}$ be the character
$\det\nolimits_{2}:P\longrightarrow\mathbb{G}_{m},\quad\begin{bmatrix}A_{1}&*\\ 0&A_{2}\end{bmatrix}\longmapsto\det(A_{2}).$
One checks that $\mathcal{O}(1)=F(E)\times_{P}k_{\det_{2}}$, $\det_{2}=\lambda_{n-r+1}+\cdots+\lambda_{n}$ in $X(T)$ and $I=(i_{1},\dots,i_{r})\in W/W_{P}$ acts on $\det_{2}$ by $I\det_{2}=\lambda_{i_{1}}+\cdots+\lambda_{i_{r}}$.
Let $h_{\mathcal{O}(1)}$ be the height function on $\operatorname{Gr}_{r}(E_{K})$ induced by the model $(\operatorname{Gr}_{r}(E),\mathcal{O}(1))$. Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest Weyl element. By Theorem A, the essential minimum of $h_{\mathcal{O}(1)}$ is $\zeta_{\operatorname{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det_{2}\rangle=\mu_{1}+\cdots+\mu_{r}$.
The only element in $\Delta\backslash\Delta_{P}$ is $\alpha=\lambda_{n-r}-\lambda_{n-r+1}$ and the function $\langle\alpha^{\vee},\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\langle\alpha^{\vee},\det_{2}\rangle<0$ and $f\longmapsto 0$. $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$. By Theorem F, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$. One sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
1.5. Acknowledgements
We thank Xinyi Yuan and Ye Tian for their consistent supports. We thank Jie Liu, Marc Besson, Hongsheng Hu, Mihai Fulger and Brian Lehmann for discussions and email correspondences.
The idea of comparing height filtrations and base loci was proposed by Huayi Chen in a private conversation. The idea of computing concave transform was inspired from discussions with François Ballaÿ. The idea of proving Proposition 3.4 by Weyl symmetry was inspired from discussions with Tao Gui. We wish to thank them especially.
2. Height filtrations and successive minima
Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$. Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The main result in this section is the following theorem:
Theorem 2.1.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
Corollary 2.2.
The filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite, and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$.
Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points $\zeta_{w}$ the successive minima. Note that our definition of successive minima are slightly different with Zhang [25]. Let $d=\dim G/P$, Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In the following, $e_{i}$ will be called Zhang’s successive minima to avoid ambiguity.
For $w\in W_{Q}\backslash W/W_{P}$, let
$\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$
be the length function, which is nothing but the dimension of the cell $C_{w}$.
Corollary 2.3.
Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)\geq d-i+1\big{\}}$.
2.1. Vector bundles
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
There exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
  1. (1)
    $E_{i}/E_{i-1}$ is semistable;
  2. (2)
    $\mu(E_{i}/E_{i-1})>\mu(E_{i+1}/E_{i})$.
This filtration is called the Harder-Narasimhan filtration of $E$. $\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$. $\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
2.2. Principal bundles
For any linear algebraic group $\Gamma/k$, let $X(\Gamma)=\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$ denote the character group of $\Gamma$. For a cocharacter $f\in X(\Gamma)^{\vee}$ and a character $\lambda\in X(\Gamma)$, we shall denote the pairing by $\langle f,\lambda\rangle$. For any $\lambda\in X(\Gamma)$, denote by $k_{\lambda}$ the one-dimensional representation on the vector space $k$ with $\Gamma$ acting by $\lambda$.
A principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
$F\times_{C}(C\times\Gamma)\longrightarrow F\times_{C}F,\quad(f,(x,g))\longmapsto(f,fg)$
is an isomorphism.
Attached to a principal $\Gamma$-bundle $F$, one has an associated cocharacter
${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism.
By the universal property of the quotient $F/H$, the assignment to a section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
2.3. Reductive groups, characters and cocharacters
Let $G$ be a connected reductive group over $k$. Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$. Let $W$ be the Weyl group and $\Delta$ be the set of simple roots with respect to $(G,B,T)$. For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
We shall consider only parabolic subgroups containing $B$. For such a parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$. Note that the natural inclusion
$X(P)\longrightarrow X(L_{P})\longrightarrow X(Z(L_{P}))$
becomes an isomorphism after tensoring with $\mathbb{Q}$. Thus we have
$X(T)_{\mathbb{Q}}\longrightarrow X(Z(L_{P}))_{\mathbb{Q}}=X(P)_{\mathbb{Q}}$
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [23, §2.1.3]. In other words, a cocharacter on $X(P)_{\mathbb{Q}}$ can be extended canonically to $X(T)_{\mathbb{Q}}$.
2.4. The canonical reduction to a parabolic subgroup
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $F_{P}$ be a reduction of $F$ to a parabolic subgoup $P$. Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter.
The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration). Similarly, among all reduction to parabolic subgroups, there is a canonical one. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
  1. (1)
    The the principal $L_{P}$ bundle $F_{P}\times_{P}L_{P}$ is semistable.
  2. (2)
    For any non-trivial character $\lambda$ of $P$ which is non-negative linear combination of simple roots, $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle>0$.
Behrend in [2] proved canonical reduction exists uniquely. Note that when $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
2.5. Arakelov geometry on flag varieties over function fields
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$.
${X=(F/P)_{K},L\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P,\mathcal{L}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$, $L=\mathcal{L}_{\lambda,K}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R},x\longmapsto\frac{\overline{\{x\}}\cdot\mathcal{L}_{\lambda}}{{\mathrm{deg}}(x)}$ be the induced height function. Here $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Let $\zeta_{\operatorname{ess}}(X):=\inf\big{\{}t:Z_{t}=X\big{\}}$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X$.
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
Note that for any $w^{\prime}\in W_{Q}$ and $\lambda\in X(T)$, we have $w^{\prime}\lambda-\lambda\in\mathbb{Z}[\Delta_{Q}]$ consequently $\langle{\mathrm{deg}}(F_{Q}),w^{\prime}\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle$. Note also that for any $\lambda\in X(P)$ and $w\in W_{P}$, $w\lambda=\lambda$. So the number $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is well-defined for any $\lambda\in X(P)$ and $w\in W_{Q}\backslash W/W_{P}$.
$\mathcal{X}_{w}$ is a closed subscheme of $\mathcal{X}$, so $\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$ induces a height function $h_{\mathcal{L}_{\lambda}}:X_{w}(\overline{K})\longrightarrow\mathbb{R}$, which is nothing but the restriction of $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ to $X_{w}(\overline{K})$. Let $\zeta_{\operatorname{ess}}(X_{w})$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$.
Theorem 2.1 follows from the following two propositions.
Proposition 2.4.
$h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$ for any $x\in C_{w}(\overline{K})$.
Proposition 2.5.
$\zeta_{\operatorname{ess}}(X_{w})=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Theorem 2.1.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
$Z_{t}=X\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}C_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t}C_{w}.$
Proof.
We write $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and $h=h_{\mathcal{L}_{\lambda}}$ to simplify notations. For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y(\overline{K}):h(y)<t\big{\}}$. By Proposition 2.4, $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}$, and we claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$. Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. But on $X_{w^{\prime}}$, Proposition 2.5 says the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$. Contradiction! Thus on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
Let $w_{0}\in W_{Q}\backslash W/W_{P}$ be the longest element. On $X=X_{w_{0}}$, when $t>\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}$ and when $t\leq\zeta_{w_{0}}$, $Z_{t}=X_{w_{0}}\backslash C_{w_{0}}=\bigcup_{w^{\prime}}X_{w^{\prime}}$ with $w^{\prime}<w_{0}$ and $\ell(w^{\prime})=\ell(w_{0})-1$. Repeating this procedure on each smaller $X_{w^{\prime}}$, we can complete the proof by induction. ∎
2.6. Proof of Proposition 2.4
For $x\in X(K)$, let $\sigma_{x}:C\longrightarrow\mathcal{X}$ be the section induced by $x:\operatorname{Spec}(K)\longrightarrow X$ via valuative criterion, and let $F_{P,x}=\sigma_{x}^{*}F$ be the corresponding reduction to $P$.
Lemma 2.6.
For $x\in X(K)$, $h_{\mathcal{L}_{\lambda}}(x)=\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$.
Proof.
By definition, $h_{\mathcal{L}_{\lambda}}(x)$ is the degree of $\sigma_{x}^{*}(\mathcal{L}_{\lambda})$ and $\langle{\mathrm{deg}}(F_{P,x}),\lambda\rangle$ is the degree of $F_{P,x}\times_{P}k_{\lambda}$. We have equality of line bundles
$\sigma_{x}^{*}(\mathcal{L}_{\lambda})=\sigma_{x}^{*}(F\times_{P}k_{\lambda})=(\sigma_{x}^{*}F)\times_{P}k_{\lambda}=F_{P,x}\times_{P}k_{\lambda}$
and the lemma follows by taking degree. ∎
Definition 2.7 (relative position).
A reduction $F_{P}$ of $F$ to $P$ is called in relative position $w\in W_{Q}\backslash W/W_{P}$ with respect to $F_{Q}$ if the image of the natural map
$F_{Q}\times_{C}F_{P}\longrightarrow G_{C},\quad(a,b)\longmapsto a^{-1}b$
lies in $QwP_{C}$. Here we implicitly apply the injections
$F_{P}\hookrightarrow F_{P}\times_{P}G\simeq F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\simeq F.$
Note that by the universal property of the quotient stacks $[G//Q\times P]$ and $[QwP//Q\times P]$, this definition coincides with the one in [23, §4.1].
Lemma 2.8.
When $x\in C_{w}(K)$, $F_{P,x}$ is of relative position $w$ with respect to $F_{Q}$.
Proof.
Note that the injection $F_{P,x}=\sigma_{x}^{*}F\longrightarrow F$ has image in $F_{Q}\times_{Q}QwP$ by the commutativity of the following diagram
Now the lemma follows since $F_{Q}\times_{C}F_{P,x}\longrightarrow C\times G$ factors as
$F_{Q}\times_{C}F_{P,x}\longrightarrow F_{Q}\times_{C}\Big{(}F_{Q}\times_{Q}QwP\Big{)}\longrightarrow$
$\Big{(}F_{Q}\times_{C}F_{Q}\Big{)}\times_{Q}QwP\longrightarrow(C\times Q)\times_{Q}QwP=C\times QwP.$
Proposition 2.9.
If $F_{P}$ is in relative position $w$ with respect to $F_{Q}$, then $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\geq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for any anti-dominant character $\lambda$.
Proof.
This is written in the proof of [23, Theorem 4.1]. See [23, §4.5.3]. ∎
Proof of Proposition 2.4.
Let $L$ be any finite field extension of $K$. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
[…]
where $C_{L}$ is the normalization of $C$ in $L$ and $F_{C_{L}}=F\times_{C}C_{L}$. Let $\mathcal{L}_{C_{L},\lambda}$ be the pullback of $\mathcal{L}_{\lambda}$ via the map $F_{C_{L}}/P\longrightarrow F/P$, which is also the line bundle associated to the character $\lambda$ of $P$ via the principal bundle $F_{C_{L}}$ on $C_{L}$. Then by Lemma 2.6,
$h_{\mathcal{L}_{\lambda}}(x)=\tfrac{1}{[L:K]}{\mathrm{deg}}(\sigma_{x}^{*}\mathcal{L}_{\lambda})=\tfrac{1}{[L:K]}{\mathrm{deg}}(\sigma_{\widetilde{x}}^{*}\mathcal{L}_{C_{L},\lambda})=\tfrac{1}{[L:K]}\langle{\mathrm{deg}}(F_{C_{L},P,\widetilde{x}}),\lambda\rangle$
where $F_{C_{L},P,\widetilde{x}}$ is the reduction of $F_{C_{L}}$ to $P$ corresponding to the rational point $\widetilde{x}\in X_{L}(L)$.
Note that the canonical reduction of $F_{C_{L}}$ is $F_{Q,C_{L}}=F_{Q}\times_{C}C_{L}$ and ${\mathrm{deg}}(F_{Q,C_{L}})=[L:K]{\mathrm{deg}}(F_{Q})$ in $X(T)_{\mathbb{Q}}^{\vee}$. By Lemma 2.8, $F_{Q,C_{L}}$ is of relative position $w$ with respect to $F_{C_{L},P,\widetilde{x}}$. We conclude by applying Proposition 2.9 to $F_{C_{L}}$. ∎
2.7. Proof of Proposition 2.5
In this subsection, we employ Ballaÿ’s theorem [1, Theorem 1.2] to compute the essential minimum. Let $\pi:F/P\longrightarrow C$ be the structure morphism. We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
Lemma 2.10 ([14], Part I, §5.18).
On $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P\subseteq\mathcal{X}$, we have $\pi_{*}\big{(}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}\big{)}=F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$.
Schieder computed the Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ in [23, §5.1]. To state his result, we fix the following notations which will be used frequently.
Definition 2.11.
Let $V$ be a finite dimensional $Q$-representation over $k$ with weight decomposition $V=\bigoplus_{\mu\in X(T)}V[\mu]$. The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}V[\mu].$
Proposition 2.12.
The Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ is $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})_{\bullet,{\mathrm{deg}}(F_{Q})}$. Moreover, the maximal slope is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Proof.
The first assertion is a slight generalization of [23, Proposition 5.1], whose original texts only consider $G$-representations restricted to $Q$, but the proof actually works for arbitrary $Q$-representations. For the second assertion, we only need to show the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$.
When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [14, Chapter 14]. For general $Q$, let $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. Note that
$\overline{QwP}/P=\bigcup_{w^{\prime}\in\pi^{-1}(w)}\overline{Bw^{\prime}P}/P.$
Since the restriction map
$H^{0}(\overline{QwP}/P,M_{\lambda})\longrightarrow H^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is surjective for each $w^{\prime}\in\pi^{-1}(w)$ and the diagonal map
$H^{0}(\overline{QwP}/P,M_{\lambda})\longrightarrow\prod_{w^{\prime}\in\pi^{-1}(w)}H^{0}(\overline{Bw^{\prime}P}/P,M_{\lambda})$
is injective, the highest weights in $H^{0}(\overline{QwP}/P,M_{\lambda})$ belong to $w\lambda+\mathbb{Z}[\Delta_{Q}]$. ∎
Proof of Proposition 2.5.
By Ballaÿ’s theorem [1, Theorem 1.2],
$\quad\quad\quad\zeta_{\operatorname{ess}}(X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
$=\lim_{n\rightarrow\infty}\frac{n\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle}{n}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
3. String polytopes and the Boucksom-Chen concave transform
Keep the notations in §2. Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. Then for the antidominant weight $\lambda:P\longrightarrow\mathbb{G}_{m}$, we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$, which is a polytope of dimension $d=\dim G/P$. Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map.
The main theorem of this section is the following:
Theorem 3.1.
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ (with respect to a specific valuation) is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. As a direct corollary, we have by [6] that
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
The integration can be computed via Weyl symmetry:
Theorem 3.2.
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$ by Theorem 2.1. Thus this theorem builds an equality between the variety height and a weighted average of the successve mimina. One may recognize this as a refinement of Zhang’s inequality of successive minima.
3.1. Okounkov bodies and the Boucksom-Chen concave transform
Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety of dimension $d$ over $\mathbf{k}$. Let $L$ be a big line bundle on $Y$, $V_{n}=H^{0}(Y,nL)$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$. Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{r}$ be a valuation with one dimensional leaves, i.e.
  • for $a\in\mathbf{k}^{\times}$ and $s\in V_{\bullet}\backslash\{0\}$, we have $\nu(as)=\nu(s)$.
  • for $s_{1},s_{2}\in V_{\bullet}\backslash\{0\}$, we have $\nu(s_{1}s_{2})=\nu(s_{1})+\nu(s_{2})$ and $\nu(s_{1}+s_{2})\geq\min\big{\{}\nu(s_{1}),\nu(s_{2})\big{\}}$.
  • the quotient space $\big{\{}s\in H^{0}(Y,nL):\nu(s)\geq a\big{\}}\Big{/}\big{\{}s\in H^{0}(Y,nL):\nu(s)>a\big{\}}$ is at most one dimensional.
Let $W_{\bullet}\subseteq V_{\bullet}$ be a graded subalgebra. We define $\Delta_{\nu}(W_{\bullet})\subseteq\mathbb{R}^{r}$ to be the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(W_{n}\backslash\{0\})$. In particular, we define the Okounkov body $\Delta_{\nu}(Y,L):=\Delta_{\nu}(V_{\bullet})$.
A fundamental theorem is $\operatorname{Vol}(L)=d!\operatorname{Vol}(\Delta_{\nu}(Y,L))$ [16, Theorem 1.15], where the left hand side is the volume of the line bundle $L$ and the right hand side is the volume of the polytope $\Delta_{\nu}(Y,L)$.
Let $\big{\{}V_{\bullet}^{t}:t\in\mathbb{R}\big{\}}$ be a decreasing $\mathbb{R}$-filtration of $V_{\bullet}$ that is multiplicative, i.e. $V_{n_{1}}^{t_{1}}\cdot V_{n_{2}}^{t_{2}}\subseteq V_{n_{1}+n_{2}}^{t_{1}+t_{2}}$. The Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$ is the function $G:\Delta_{\nu}(V_{\bullet})\longrightarrow\mathbb{R}\cup\{-\infty\}$, $x\longmapsto\sup\big{\{}t\in\mathbb{R}:x\in\Delta_{\nu}(V_{\bullet}^{t})\big{\}}$ [6]. $G$ is concave, as indicated in the name. In particular, $G$ is continuous in the interior.
3.2. String polytopes and weight polytopes
Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. In [19, 3], the authors constructed the string valuation $\nu=\nu_{\underline{w_{0}}}:H^{0}(G/P,nM_{\lambda})\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$. The image $\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})=\big{\{}v_{1},\dots,v_{l}\big{\}}$ is in bijection (the string parametrization) with the so-called dual crystal basis $\big{\{}e_{v_{1}}^{(n\lambda)},\cdots,e_{v_{l}}^{(n\lambda)}\big{\}}$ of $H^{0}(G/P,nM_{\lambda})$ (see [16, §3] for a detailed exposition). Moreover, the dual crystal basis is compatible with the weight decomposition
$H^{0}(G/P,nM_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,nM_{\lambda})[\mu],$
that is, for each $i=1,\dots,l$, $e_{v_{i}}^{(n\lambda)}\in H^{0}(G/P,nM_{\lambda})[\mu]$ for some $\mu$.
The string polytope $\Delta_{\lambda}$ is the Okounkov body $\Delta_{\nu}(G/P,M_{\lambda})$, i.e. it is the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})$. Note that $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ is in bijection with the dual crystal basis of $H^{0}(G/P,nM_{\lambda})$.
The weight polytope $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the convex hull of the Weyl orbit $W\lambda$ in $X(T)_{\mathbb{R}}$. Note that $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ is in bijection with weights apprearing in $H^{0}(G/P,nM_{\lambda})$. By taking weights, one has weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ which is affine and sends $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ to $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$
3.3. Proof of Theorem 3.1
Let $V_{n}=H^{0}(G/P,nM_{\lambda})$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$. For $t\in\mathbb{R}$, let $V_{n}^{t}=H^{0}(G/P,nM_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$ as in Proposition 2.12, and $V_{\bullet}^{t}=\bigoplus_{n\geq 0}V_{n}^{nt}$. Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$ be the string valuation. Let $\varphi:\Delta_{\nu}(G/P,M_{\lambda})=\Delta_{\lambda}\longrightarrow\mathbb{R}$ be the Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$.
Theorem 3.3.
$\varphi(x)=\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
Proof.
$V_{n}^{nt}=\bigoplus_{\mu}V_{n}[n\mu]$ for $\mu\in\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ with $\langle{\mathrm{deg}}(F_{Q}),n\mu\rangle\geq nt$, and hence $V_{n}^{nt}=\bigoplus_{x}k\cdot e_{nx}^{(n\lambda)}$ for $x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ with $\langle{\mathrm{deg}}(F_{Q}),p(nx)\rangle\geq nt$. Therefore by definition,
$\quad\quad\quad\Delta(V_{\bullet}^{t})$ $=\text{closure of }\bigcup\nolimits_{n}\big{\{}x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
=$=\big{\{}x\in\Delta_{\lambda}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
which concludes the proof. ∎
Consider the graded linear series $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}H^{0}(X,nL)$. Fix an isomorphism $F_{Q,K}\simeq Q_{K}$. Then
(3.1) $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}\simeq\bigoplus_{n\geq 1}H^{0}(G/P,nM_{\lambda})_{K}.$
Consider the filtrations of $\widetilde{V}_{n}=\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}$ induced by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})$ as in Proposition 2.12
$\widetilde{V}_{\bullet}^{t}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}\simeq\bigoplus_{n\geq 0}\Big{(}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}.$
Via the string parametrization of the dual crystal basis, the string valuation $\nu$ on $H^{0}(G/P,nM_{\lambda})$ extends canonically to a valuation $\widetilde{\nu}$ on $H^{0}(G/P,nM_{\lambda})_{K}$. This induces a valuation on $\widetilde{V}_{n}\simeq H^{0}(G/P,nM_{\lambda})_{K}$ via the isomorphism (3.1). We shall abuse notation and also call it $\widetilde{\nu}$.
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ with respect to $\widetilde{\nu}$ is the one associated to $\left(\widetilde{V}_{\bullet},\widetilde{\nu},\big{\{}\widetilde{V}_{\bullet}^{t}\big{\}}\right)$, and obviously it coincides with the one associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$. Theorem 3.1 follows.
3.4. Proof of Theorem 3.2
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. By [6], we have
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
Proposition 3.4.
For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
$\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}f(w\lambda).$
Proof of Theorem 3.2.
Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.4. ∎
The rest of this subsection is devoted to the proof of Proposition 3.4. We start with two lemmata.
Lemma 3.5.
The pushforward measure $p_{*}\mu$ on $\mathcal{P}_{\lambda}$ is $W$-invariant.
Proof.
Note that the measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measures $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e. $\mu=\lim_{n}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$. Consider the map $\Delta_{n\lambda}\longrightarrow\mathcal{P}_{n\lambda}$ and note that it sends integral points to its corresponding weights. Dividing by n, we see that
$p_{*}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}=\operatorname{mult}_{n}\delta_{\mathcal{P}_{\lambda}\cap\frac{1}{n}X(T)},$
where $\operatorname{mult}_{n}$ is the multiplicity function
$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,M_{n\lambda})[n\mu].$
Since both $\operatorname{mult}_{n}$ and $\delta_{\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)}$ are $W$-invariant, we deduce that
$p_{*}\mu=\lim_{n}p_{*}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$
is also $W$-invariant. ∎
Lemma 3.6.
The map $\mathcal{P}_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$, $\mu\longmapsto\sum_{w\in W}w\mu$ is constant.
Proof.
We claim that for any $\mu_{1},\mu_{2}\in X(T)$ such that $\mu_{1}|_{Z(G)}=\mu_{2}|_{Z(G)}$,
$\sum_{w\in W}w\mu_{1}=\sum_{w\in W}w\mu_{2}.$
In fact, for any simply connected cover $G^{sc}\longrightarrow G$, the preimage $T^{sc}$ of $T$ is a maximal torus in $G^{sc}$ and $W=W(G,T)=W(G^{sc},T^{sc})$. Via the natural map $X(T)\longrightarrow X(T^{sc})$, we are reduced to the case $G$ is simply connected. In this case, $T^{W}=Z(G)$ and the desired equality follows directly.
$H^{0}(G/P,M_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,M_{\lambda})$ share a same restriction to $Z(G)$. So by the above claim, the lemma is true for lattice points $\mathcal{P}_{\lambda}\cap X(T)$. Note that the map $\mu\longmapsto\sum_{w\in W}w\mu$ is linear, we are done. ∎
Proof of Proposition 3.4.
Let $\mathcal{P}_{\lambda,+}$ be the intersection of $\mathcal{P}_{\lambda}$ with the anti-dominant cone $\big{\{}\langle\alpha^{\vee},\lambda\rangle<0:\alpha\in\Delta\backslash\Delta_{P}\big{\}}$. Note that the subgroup $W_{P}\subseteq W$ acts trivially on $\mathcal{P}_{\lambda}$. By Lemma 3.5, one has
$\int_{\Delta_{\lambda}}f\circ p\;\mathrm{d}\mu=\int_{\mathcal{P}_{\lambda}}f\;\mathrm{d}p_{*}\mu=\int_{\mathcal{P}_{\lambda,+}}\Big{(}\sum_{w\in W/W_{P}}f(wx)\Big{)}\;\mathrm{d}p_{*}\mu.$
By the linearity of $f$ and Lemma 3.6, it moreover equals to
$\int_{\mathcal{P}_{\lambda,+}}f\Big{(}\sum_{w\in W/W_{P}}wx\Big{)}\;\mathrm{d}p_{*}\mu=\Big{(}\sum_{w\in W/W_{P}}f(w\lambda)\Big{)}\int_{\mathcal{P}_{\lambda,+}}1\;\mathrm{d}p_{*}\mu.$
On the other hand,
$\operatorname{Vol}(\Delta_{\lambda})=\int_{\mathcal{P}_{\lambda}}1\;\mathrm{d}p_{*}\mu=\#{W/W_{P}}\int_{\mathcal{P}_{\lambda,+}}1\;\mathrm{d}p_{*}\mu.$
Taking ratios, we obtain the desired equality. ∎
We end this section by the case $\operatorname{Gr}(2,4)$.
Example 3.7.
Let $G=\operatorname{GL}_{4}$, $E$ be a vector bundle on $C$ of rank $4$ and consider the $\operatorname{Gr}(2,4)$-bundle $\operatorname{Gr}_{2}(E)$. In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [15, Remark 2.4]. Identify $X(T)_{\mathbb{Q}}$ with $\bigoplus_{i}\mathbb{Q}\lambda_{i}$ as before. The line bundle $\mathcal{O}(1)$ corresponds to the character $\lambda=(0,0,1,1)$.
The Gelfand-Zetlin polytope $\operatorname{GZ}_{\lambda}$ is defined by $0\leq x_{2}\leq x_{1}\leq x_{3}\leq 1$ and $x_{2}\leq x_{4}\leq x_{3}$ in $\mathbb{R}^{4}$, with vertices $(0,0,0,0)$, $(0,0,1,0)$, $(0,0,1,1)$, $(1,0,1,0)$, $(1,0,1,1)$ and $(1,1,1,1)$. The weight polytope $\mathcal{P}_{\lambda}$ is the convex hull of $(0,0,1,1)$, $(0,1,0,1)$, $(0,1,1,0)$, $(1,1,0,0)$, $(1,0,1,0)$ and $(1,0,0,1)$, the weight map $p:\operatorname{GZ}_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ is
$(x_{1},x_{2},x_{3},x_{4})\longmapsto(x_{1},-x_{1}+x_{2}+x_{3},-x_{2}-x_{3}+x_{4},-x_{4})+(0,0,1,1)$
and ${\mathrm{deg}}(F_{Q})$ is the function on $X(T)_{\mathbb{Q}}$ given by $\lambda_{i}\longmapsto\mu_{i}$ as in §1.4.
A direct computation shows that
$\operatorname{Vol}(\operatorname{GZ}_{\lambda})=\frac{1}{12},\quad\int_{\operatorname{GZ}_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu=\frac{1}{24}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4}).$
Thus $h(X)=\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})$.
On the other hand, $W/W_{P}=S_{4}\big{/}(S_{2}\times S_{2})$, #$\#W/W_{P}=6$ and
$\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\sum_{1\leq i<j\leq 4}\mu_{i}+\mu_{j}.$
Thus $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})=h(X)$, which is predicted by Theorem 3.2.
4. Base loci and movable cones
Keep the notations in §2. We fix a closed point $p_{0}\in C$ and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$. For any $w\in W_{Q}\backslash W/W_{P}$, let $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$. Let $t\in\mathbb{Q}$. The main theorem of this section is a computation of base loci.
Theorem 4.1.
Let $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
Comparing Theorem 2.1 and Theorem 4.1, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ agrees with the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
As a corollary, we compute the movable cones on flag bundles. (Please see §4.3 for the notations.)
Theorem 4.2.
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
In particular, let $\mathcal{L}$ be a line bundle on $F/P$ and $e_{k}(h_{\mathcal{L}})$ be the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$, we have
Corollary 4.3.
$\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$.
4.1. Augmented base loci and the restricted volume
Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety over $\mathbf{k}$. Let $L$ be a line bundle on $Y$ and $V\subseteq H^{0}(Y,L)$ be a subspace. We shall denote by $\mathrm{Bs}(V)$ the base locus of $V$.
The stable base locus of $L$ is defined as $\mathrm{B}(L)=\bigcap_{n\geq 1}\mathrm{Bs}(H^{0}(Y,nL))$. The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$ for any ample line bundle $A$. This is independent of the choice of $A$.
Let $Z$ be a closed subvariety of $Y$ of dimension $d$. The restricted volume is defined as
$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}\Big{(}H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z})\Big{)}}{n^{d}/d!}.$
The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for readers’ convenience.
Theorem 4.4 ([11], Theorem C).
$\mathrm{B}_{+}(L)$ is the union of all closed subvarieties $Z$ with $\operatorname{Vol}(Y|Z,L)=0$.
Let $N^{1}(Y)$ be the Néron-Severi group of $Y$. It is proved in [10] that augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $\mathcal{Y}\longrightarrow C$ be a projective flat family over the curve $C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$. For any scheme theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$. Let $\mathcal{Z}$ be a closed subvariety of $\mathcal{Y}$, we denote by $\mathcal{Z}_{p}$ the fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$. We denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
Lemma 4.5.
$\mathrm{B}(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}\left(\mathrm{Im}\Big{(}H^{0}(\mathcal{Y},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p})\Big{)}\right)$.
4.2. Proof of Theorem 4.1
Let $M_{\lambda}=G\times_{P}k_{\lambda}$, which is an ample line bundle on $G/P$. Recall that we have the decomposition
$H^{0}(G/P,M_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,M_{\lambda})[\mu]$
and the filtration
$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\mu].$
We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\mu]$.
Lemma 4.6.
Take $w\in W/W_{P}$. Then
  1. (1)
    $s\neq 0\in H^{0}(G/P,M_{\lambda})[w\lambda]$Longrightarrow $s(x)\neq 0$ for any $x\in BwP/P$.
  2. (2)
    $s\in H^{0}(G/P,M_{\lambda})[\mu]$ for $\mu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
Proof.
Consider the restriction map
$\mathrm{Res}:\ H^{0}(G/P,M_{\lambda})\longrightarrow H^{0}(\overline{BwP}/P,M_{\lambda}),$
which is $B$-equivariant and surjective. Since $H^{0}(G/P,M_{\lambda})[w\lambda]$ and $H^{0}(\overline{BwP}/P,M_{\lambda})[w\lambda]$ are both one-dimensional, one has $\mathrm{Res}(s)\neq 0$ for any $s\neq 0\in H^{0}(G/P,M_{\lambda})[w\lambda]$.
By the closedness of the vanishing locus of $\mathrm{Res}(s)$, there exists $x\in BwP/P$ such that $\mathrm{Res}(s)(x)\neq 0$. Note that by the highest weight theory, $B$ acts on $\mathrm{Res}(s)$ via $w\lambda$ and $BwP/P$ is a single $B$-orbit under translation. Consequently, $\mathrm{Res}(s)(x)\neq 0$ for any $x\in BwP/P$.
As for $(2)$, for any $\mu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\mu]=0$. Consequently, $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\mu]$. ∎
Let $p$ be a scheme theoretical point of $C$ and consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}.$
We have
Lemma 4.7.
$\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
Proof.
Let $F_{T,\kappa(p)}$ be a reduction of $F_{Q,\kappa(p)}$ to $T$. For any $\mu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\mu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$. For $w^{\prime}\in W/W_{P}$, set $C_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}Bw^{\prime}P/P$ and $X_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}\overline{Bw^{\prime}P}/P$. Then $\mathcal{C}_{w,p}=\coprod_{\pi(w^{\prime})=w}C_{w^{\prime}}$ and $\mathcal{X}_{w,p}=\bigcup_{\pi(w^{\prime})=w}X_{w^{\prime}}$ where $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ is the natural projection.
Since the principal $T$-bundle $F_{T,\kappa(p)}$ is trivial, we deduce that $s(x)\neq 0$ for any $x\in C_{w^{\prime}}$, any $w^{\prime}\in\pi^{-1}(w)$, and any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[w^{\prime}\lambda]$ from Lemma 4.6(1). Thus if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t$, $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[w^{\prime}\lambda]\subseteq H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$ for any $w^{\prime}\in\pi^{-1}(w)$ and consequently, $x\not\in\operatorname{Bs}(H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\geq t})$ for any $x\in C_{w^{\prime}}$. This proves $\operatorname{Bs}(H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})})\subseteq\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
On the other hand, we claim that if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$, then for any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$, the restriction of $s$ to $X_{w}$ is zero. In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]$ for some $\mu$ and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t$ implies that $\mu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\mu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$). By Lemma 4.6(2), the restriction of $s$ to $X_{w^{\prime}}$ is zero for each $w^{\prime}$. We win. ∎
Theorem 4.1.
Let $B_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}.$
Proof.
We may assume that $t\in\mathbb{N}$ by taking a sufficiently large tensor power of $\mathcal{L}_{\lambda}-tf$. Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle:H^{0}(G/P,M_{\lambda})[\mu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\mu\rangle>t\}$. Take $E=F_{Q}\times_{Q}H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}\otimes\mathcal{O}_{C}(-tx)$, which is a subbundle of $\pi_{*}(\mathcal{L}_{\lambda}-tf).$ Moreover, $\mu_{\min}(E)=t_{0}-t>0$, i.e. $E$ is ample.
For each scheme-theoretical point $p\in C$, we have $E(p):=E\otimes_{C}\kappa(p)=F_{Q,\kappa(p)}\times H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}.$ The ampleness of $E$ gives that
$H^{0}(C,\mathrm{Sym}^{m}E)\otimes_{k}\kappa(p)\longrightarrow\mathrm{Sym}^{m}(E(p))$
is surjective for every $m\gg 0$. Since $\pi_{*}(\mathcal{L})\otimes_{C}\kappa(p)=H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$, we have the commutative diagram
which implies that $V:=\mathrm{Im}(\mathrm{Sym}^{m}E(p)\rightarrow H^{0}(\mathcal{X}_{p},m\mathcal{L}_{\lambda,p}))$ is contained in the image $\mathrm{Im}(H^{0}(C,\pi_{*}(m\mathcal{L}_{\lambda}-tmf))\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{X}_{p},m\mathcal{L}_{\lambda,p}))=\mathrm{Im}(H^{0}(\mathcal{X},m\mathcal{L}_{\lambda}-tmf)\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{X},m\mathcal{L}_{\lambda,p})).$ Hence
$\mathrm{B}(\mathcal{L}_{\lambda}-tf)_{p}\subseteq\mathrm{Bs}(V)=\mathrm{Bs}(E(p))=\mathcal{X}_{p}\backslash\coprod\nolimits_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w,p}$
by applying Lemma 4.7 to $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t_{0},{\mathrm{deg}}(F_{Q})}$. Consequently,
(4.1) $\mathrm{B}(\mathcal{L}_{\lambda}-tf)\subseteq\mathcal{X}\backslash\coprod\nolimits_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle-t^{\prime}>0$ for any $\mu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\mu]\not=0.$
Take sufficiently large $n$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda^{\prime}\rangle-t^{\prime}<n(t_{0}-t)$ for any $w\in W_{Q}\backslash W/W_{P}$. Thus $\langle{\mathrm{deg}}(F_{Q}),w(n\lambda-\lambda^{\prime})\rangle>nt-t^{\prime}$ if and only if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t_{0}.$ Similarly as in (4.1), we obtain that
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)\subset\mathrm{B}(n(\mathcal{L}_{\lambda}-tf)-\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f)\subseteq\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
On the other hand, for any $w\in W_{Q}\backslash W/W_{P}$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t$, since
$\limsup_{m\rightarrow+\infty}\frac{\mu_{\max}(\pi_{*}(m(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{w}})}{n}=\zeta_{1}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=\zeta_{w}-t\leq 0,$
we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [9, Theorem 2.4] deduced from the Riemann-Roch theorem. Therefore $\operatorname{Vol}(\mathcal{X}|\mathcal{X}_{w},\mathcal{L}_{\lambda}-tf)=0$, which implies that $\mathcal{X}_{w}\subseteq B_{+}(\mathcal{L}_{\lambda}-tf)$ due to Theorem 4.4. In conclusion, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
4.3. Movable cones and proof of Theorem 4.2
Let $Y$ be a projective smooth variety over $\mathbf{k}$. Recall
  • The big cone $\mathrm{Big}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of big divisors. Its closure is called Pseudo-effective cone, denoted as $\mathrm{Psef}(Y)$.
  • The ample cone $\mathrm{Amp}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is the cone of numerical classes of ample divisors. Its closure is called nef cone, denoted as $\mathrm{Nef}(Y)$.
  • For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is a open in $N^{1}(Y)_{\mathbb{R}}$.
We have obvious inclusions $\operatorname{Mov}^{d}(Y)\subseteq\cdots\subseteq\operatorname{Mov}^{1}(Y)$. Note that $\operatorname{Mov}^{1}(Y)$ is just the big cone and $\operatorname{Mov}^{d}(Y)$ is just the ample cone.
Now we go back to the setting of §2. Denote by
$c:X(P)\longrightarrow\mathrm{Pic}(G/P),\quad\lambda\longmapsto M_{\lambda}$
the character map. We have an exact sequence [20, Theorem 18.32]:
$0\longrightarrow X(G)\longrightarrow X(P)\longrightarrow\mathrm{Pic}(G/P)\longrightarrow\mathrm{Pic}(G)\longrightarrow\mathrm{Pic}(P)\longrightarrow 0$
Since $G$ and $P$ are connected linear algebraic groups over a characteristic zero field, both $\mathrm{Pic}(G)$ and $\operatorname{Pic}(P)$ are finite. Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since it is smooth and Fano. Therefore the linear map $X(P)_{\mathbb{R}}\rightarrow N^{1}(G/P)_{\mathbb{R}}$ is surjective. It is well-known that $M_{\lambda}$ is ample $\Longleftrightarrow$ $M_{\lambda}$ is big $\Longleftrightarrow$ $\lambda$ is strictly anti-dominant, thus
$\mathrm{Amp}(G/P)=\mathrm{Big}(G/P)=\big{\{}c(\lambda):\langle\alpha^{\vee},\lambda\rangle<0\text{ for all $\alpha\in\Delta\backslash\Delta_{P}$}\big{\}}$
[…]
Then we compute $\mathrm{Mov}^{i}(F/P)$.
Lemma 4.8.
Let $f\in N^{1}(F/P)$ be the class of a vertical fiber. Then every element in $N^{1}(F/P)$ can be writen as $\mathcal{L}_{\lambda}-tf$.
Proof.
Apply [20, Theorem 18.32] to the $G$-bundle $F\longrightarrow C$. We get an exact sequence
$X(P)\longrightarrow\operatorname{Pic}(F/P)\longrightarrow\operatorname{Pic}(F)\longrightarrow\operatorname{Pic}(P).$
Apply it to the $P$-bundle $F\longrightarrow F/P$. We get another exact sequence
$X(G)\longrightarrow\operatorname{Pic}(C)\longrightarrow\operatorname{Pic}(F)\longrightarrow\operatorname{Pic}(G).$
Note that after tensoring $\mathbb{R}$, $\operatorname{Pic}(P)_{\mathbb{R}}=\operatorname{Pic}(G)_{\mathbb{R}}=0$.
Consider the diagram
where $g$, $\phi$ and $\psi$ are just pullbacks. Note that $g$ and $\psi$ are surjective. Let $x\in\operatorname{Pic}(F/P)$ such that $g(x)=\psi(a)$ for some $a\in\operatorname{Pic}(C)_{\mathbb{R}}$. Then $g(x-\phi(a))=0$ and the first line is exact at the middle, so $x-\phi(a)=f(b)$. ∎
Thanks to the above description, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle\alpha^{\vee},\lambda\rangle$.
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\mathcal{L}_{\lambda}-tf$ to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t$.
To see they are well-defined, it suffices to deal with the case $\mathcal{L}_{\lambda}$ equals to a multiple of $f$. In this case, we see that $M_{\lambda}$ is a trivial line bundle on $G/P$ by restricting $\mathcal{L}_{\lambda}$ to a fiber. This implies $\lambda\in X(G)$ by the exact sequence $0\longrightarrow X(G)\longrightarrow X(P)\longrightarrow\operatorname{Pic}(G/P)$. Then $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}=\pi^{*}(F\times_{G}k_{\lambda})$ where $\pi$ is the structure map $F/P\longrightarrow C$. In particular, $\mathcal{L}_{\lambda}=\langle{\mathrm{deg}}(F),\lambda\rangle f$ in $N^{1}(F/P)_{\mathbb{R}}$.
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle$ sends $f$ to zero and sends $\lambda\in X(G)$ also to zero. This proves $\langle\alpha^{\vee},\cdot\rangle$ is well-defined.
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ sends $\langle{\mathrm{deg}}(F),\lambda\rangle f$ to $\langle{\mathrm{deg}}(F),\lambda\rangle$ since $\lambda\in X(G)$ is invariant under Weyl group action and sends $\lambda\in X(G)$ also to $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\langle{\mathrm{deg}}(F_{Q}),\lambda\rangle=\langle{\mathrm{deg}}(F),\lambda\rangle$ since $F\times_{G}k_{\lambda}=F_{Q}\times_{Q}G\times_{G}k_{\lambda}=F_{Q}\times_{Q}k_{\lambda}$. This proves $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ is well-defined.
Theorem 4.2.
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by
  1. (1)
    $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$, and
  2. (2)
    $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
Proof.
The sufficiency follows immediately after Theorem 4.1. For the necessity, let $\mathcal{L}_{\lambda}-tf$ be a line bundle on $F/P$ such that $\mathrm{codim}(\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf))\geq k$. Then in particular $\mathcal{L}_{\lambda}-tf$ is big. This impies $M_{\lambda}$ is big on $G/P$, which is equivalent to $\langle\alpha^{\vee},\lambda\rangle<0$ for all $\alpha\in\Delta\backslash\Delta_{P}$. Then $\lambda$ is strictly anti-dominant and Theorem 4.1 applies, saying $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t>0}\mathcal{C}_{w}$. It has $\operatorname{codim}\geq k$Longleftrightarrow $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle-t>0$ for all $w$ with $\ell(w)\geq n-k+1$. ∎